OpenStudy (anonymous):

2x - y - 4 = 0 3x + y - 9 = 0 What is the solution set of the given system? a){(6/5, 13/5)} b){(13/5, 6/5)} c){(5, 6)} d){(6, 5)}

OpenStudy (anonymous):

adding both equations we have, 5x-13=0 5x=13 x=13/5 putting this value of x in 2x-y-4=0 2(13/5)-y-4=0 26/5-4=y y=(26-20)/5 y=6/5 so {13/5,6/5} is the solution set.